# Multi-unit soft sensing permits few-shot learning in virtual flow metering

### v2

@@ Abstract
Recent literature has explored various ways to improve soft sensors by utilizing learning algorithms with transferability. A performance gain is generally attained when knowledge is transferred among strongly related soft sensor learning tasks. A particularly relevant case for transferability is when developing soft sensors of the same type for similar, but physically different processes or units. Then, the data from each unit presents a soft sensor learning task, and it is reasonable to expect strongly related tasks. Applying methods that exploit transferability in this setting leads to what we call multi-unit soft sensing.
This paper formulates multi-unit soft sensing as a probabilistic, hierarchical model, which we implement using a deep neural network. The learning capabilities of the model are studied empirically on a large-scale industrial case by developing virtual flow meters (a type of soft sensor) for 80 petroleum wells. We investigate how the model generalizes with the number of wells/units. Interestingly, we demonstrate that multi-unit models learned from data from many wells, permit few-shot learning of virtual flow meters for new wells. Surprisingly, regarding the difficulty of the tasks, few-shot learning on 1-3 data points often leads to high performance on new wells.
[…]
[SS]organization=Solution Seeker AS, city=Oslo, country=Norway
[…]
[NTNU]organization=Norwegian University of Science and Technology, city=Trondheim, country=Norway
@@ 1.1 Soft sensing
In essence, a soft sensor is a mathematical model that operates on available process measurements ($x$) to infer a variable of interest ($y$). Arguably, the most common form of a soft sensor is $\mathrm{E}\mathopen{}\left[y\,|\,x,\theta\right]\mathclose{}=F(x;\theta)$, where the conditional expected value of $y$ is modeled by a function $F$ with arguments $x$ and parameters $\theta$. Two important assumptions in the development of a soft sensor are i) that the process measurements, $x$, are informative of $y$, and ii) that the process measurements arrive with a higher frequency than measurements of $y$. When applied for online prediction, a soft sensor is used to indirectly monitor $y$ at times when only $x$ is measured. Furthermore, when $y$ is a key performance indicator, the process can be optimized based on the inferred values.
A soft sensor can be classified as being either model-driven or data-driven, based on the specification of $F$ [2]. In a model-driven soft sensor, $F$ is derived from first-principles. For example, $F$ may be a mechanistic model whose parameters represent physical properties of the process being modeled. In a data-driven soft sensor, $F$ is a generic function learned from data, e.g. a linear regression model or a neural network. Soft sensors that combine the model-driven and data-driven approaches are sometimes referred to as hybrid soft sensors. Regardless of the soft sensor type, the model should be calibrated to available measurement pairs $\mathcal{D}=\{(x,y)\}$ before being used for prediction. Commonly used methods include least-squares or maximum likelihood estimation (MLE). When a prior is available, maximum a posteriori (MAP) estimation or Bayesian inference can be used to calibrate the model [9].
@@
The disadvantages of model-driven soft sensors have motivated much of the development of data-driven soft sensors. The attraction to data-driven soft sensing comes from its promise to simplify and reduce the cost of modeling by learning the model from historical data using statistical inference or machine learning techniques. A challenge with data-driven soft sensing is to learn a model from information-poor data, where the paucity of information may be due to a low data volume, frequency, variety, or quality [2]. This challenge is closely tied to the motivation for implementing a soft sensor, which is to infer a key variable that is measured infrequently. Put differently, a soft sensor is most valuable in information-poor environments, where one would expect data-driven methods to perform poorly. To achieve a good predictive performance in such environments, it is imperative to be data efficient. Model-driven soft sensors achieve this by having few parameters and strong priors. On the other hand, data-driven soft sensors, which may have more model parameters than observations to learn from, often rely on generic regularization techniques and weak priors. In effect, this means that their model is learned from scratch for each application. Thus, we argue that data efficiency is a pressing issue in data-driven soft sensing.
@@ 1.2 Multi-unit soft sensing
One strategy to making data-driven soft sensors more data efficient is to exploit transferability [10]. A general description of transferability is the ability of a learning algorithm to transfer/share knowledge between different learning tasks to generalize better. In soft sensing, transferability entails learning across multiple soft sensor tasks. If the tasks are strongly related (i.e., share important characteristics), one may expect transferability to help soft sensors generalize. Various ways to employ transferability in soft sensing are covered in the following Section 2, where we elaborate on how different learning methods have been used to transfer knowledge between different soft sensor tasks.
@@
A particularly relevant case for transferability is when the same type of soft sensor is to be developed for several similar, but physically different processes. An example is when a product quality indicator (soft sensor) is to be developed for several, almost identical production lines. In this case, it is reasonable to expect the different soft sensor tasks to be strongly related. We call the application of transferability in this setting multi-unit soft sensing, where unit refers to a physical process being modeled by a soft sensor. [^1] In other works, a unit may be referred to as an entity, subject, object, or realization for which we have multiple observations. In statistical terms, it would be called a unit of analysis with repeated measurements. Also note that a unit may represent a unit operation, which is a term used in chemical engineering to refer to a basic step in a larger process. For example, units may refer to a set of pumps, valves, solar panels, distillation columns, geothermal wells, or robots. The setting of multi-unit soft sensing is illustrated in Fig. 1.
@@
Data-driven multi-unit soft sensing will be framed in mathematical terms in our problem statement in Section 3. Interestingly, the framing allows us to draw a parallel to model-driven soft sensing. Suppose that we have access to a large multi-unit dataset, in terms of the number of observations and units. One could then expect that a learning algorithm with transferability will converge to a prototypical representation of the data (which we can view as a strong prior model). When facing new units, this prototypical representation can be leveraged to quickly learn a soft sensor from a few data points. In machine learning terminology, this would be an example of few-shot learning, where a model performs well on a task after being calibrated to a few (often less than ten) data points [11, 12]. The parallel to model-driven soft sensing is that a first-principles model and a learned prototypical model take the same role of representing the general process.
@@
Figure 1: Illustration of multi-unit data generation. Here, each unit $F_{i}$ corresponds to a physical unit performing the same basic operation. The $M$ units are similar to each other, and can be thought of as different realizations of some “prototype” unit, denoted here by $F$. Still, each unit also has its own unique physical characteristics, resulting in datasets $\mathcal{D}_{i}$ with different distributions.
@@ 1.3 Virtual flow metering
Data-driven VFMs attempt to address the cost issue with first principles VFMs by lowering the model setup and maintenance costs. Notable examples include VFMs based on deep learning, which have shown promising results [19, 20, 21, 13]. However, it is commonly believed that their successful use requires large amounts of data [13]. Such a requirement would limit their applicability, since the wells with the worst instrumentation (and hence, the least data) are typically also the wells were VFMs are needed the most. Thus, the application of data-driven VFM to wells with few observations is a problem of high interest.
@@ 1.5 Paper outline
The paper proceeds by discussing related works in Section 2, before giving a mathematical problem formulation in Section 3. The multi-unit soft sensor model is then described in Section 4, followed by the derivation of the learning method in Section 5. An empirical study based on a large-scale industrial dataset is presented in Section 6. Finally, concluding remarks are given in Section 7. Some implementation details and additional results are gathered in the appendix.
@@ 1.6 Notation
The normal distribution is denoted by $\mathcal{N}(\mu,\Sigma)$, where $\mu$ is a mean vector and $\Sigma$ is a covariance matrix. If $X$ is a normally distributed random variable, we denote a sample by $x\sim\mathcal{N}(x\,|\,\mu,\Sigma)$. Generally, we use the letter $p$ to denote a probability distribution, $p(x)$, or a conditional probability distribution, $p(y\,|\,x)$. We sometimes use the letter $q$ for probability distributions that are approximative.
@@ 2 Related works
Finally, knowledge can be transferred between soft sensors for similar, but physically different processes. This setting is referred to as cross-entity transfer learning in [10]. The idea behind cross-entity transfer learning, is that, if two entities/processes/units are physically similar, data from one unit will be informative to the operation of the other unit. Thus, with data from multiple units, soft sensor performance may be improved by pooling the data and learning across units. This learning paradigm has high potential since many soft sensors are developed for processes that are commonly found in the industry; consider a soft sensor for a specific pump unit, which may be of a model produced in large numbers and installed in many process systems. Despite the high potential, few examples of cross-entity transfer learning can be found in the literature (one reason may be that multi-unit datasets are rarely shared). One example is given in [20], where a data-driven virtual flow meter is learned from data from many oil wells using multi-task learning. In our work, we prefer to use the term multi-unit soft sensing since it reflects that learning happens across units without specifying by which means (e.g. transfer learning or multi-task learning).
@@ 3 Problem statement
[…]
We now state that the goal of multi-unit soft sensing is to learn from $\mathcal{D}$ a distribution $q(y\,|\,x,c)$ that approximates $p(y\,|\,x,c)$. We contrast this to the single-unit soft sensing, where $p_{i}(y\,|\,x)$ is approximated using $\mathcal{D}_{i}$, for $i=1,\ldots,M$; that is, where $M$ soft sensors are learned separately.
Our work aims to study empirically and theoretically the predictive performance of the approximation $q(y\,|\,x,c)$. First, we want to investigate how performance scales with the number of units $M$. Second, we want to examine the few-shot learning performance of $q(y\,|\,x,c)$ as the context $c$ is calibrated to few observations of a new unit (whose data was not used to learn $q$).
@@ 4 Model description
We consider a hierarchical model of the observed target variables for a set of units. The model has three levels, as illustrated graphically in Fig. 2. The top-level has universal parameters ($\theta$) that are shared by all units and observations. The middle unit layer has parameters ($c$ and $\tau$) that capture the specificity of units. At the inner observation level, the target variable ($y$) is explained by the explanatory variables ($x$) and the context ($c$) of the observed unit. The target observation is subject to a measurement error ($\varepsilon$) with a unit-specific variance ($1/\tau$). The relationship between these variables is modeled using the universal parameters ($\theta$). The hierarchical structure allows the model to capture variations at the level of units and observations. A mathematical description of the model is given next.
@@
where $\varepsilon_{ij}$ is the error term. The error, here interpreted as measurement noise, is modeled as $\varepsilon_{ij}\sim\mathcal{N}(0,1/\tau_{i})$, where the precision (or variance $\sigma_{i}^{2}=1/\tau_{i}$) is allowed to differ between units. The function $f$ is modeled by a neural network with parameters $(c,\theta)$ and arguments $x$. The specific architecture of $f$ is described in Appendix A.
@@
The model in (1) results from several simplifying assumptions: 1) the dimension of the context variables, $K$, is fixed and considered a design choice; 2) the model is conditional on $x$, and thus not a generative model of the observed data; 3) the model is homoscedastic since the measurement noise variance, $\sigma_{i}^{2}$, is fixed (per unit). Assumption 2 leads to a discriminative model which is suitable when the soft sensor will be used for on-line prediction. Relaxing assumption 3, for example by allowing $\sigma_{i}$ to vary with $x$, would give a more complex heteroscedastic model, which may be harder to learn in a data-scarce setting.
@@ 4.2 Likelihood
$\ell(\bm{c},\bm{\tau},\theta\,|\,\mathcal{D}):=$ $\log p(\bm{y}\,|\,\bm{x},\bm{c},\bm{\tau},\theta)$ (3)
@@
We use the short-form notation $\ell(\bm{c},\bm{\tau},\theta\,|\,\mathcal{D})$ for the log-likelihood when it is considered as a function of the parameters ($\bm{c}$, $\bm{\tau}$, $\theta$), given data $\mathcal{D}=(\bm{x},\bm{y})$. In Section 5, we show how to perform MAP estimation of the parameters, given the log-likelihood in (3) and log-prior in (2).
@@ 5 Learning method
In model learning, the posterior density of the parameters $(\bm{c},\bm{\tau},\theta)$ is of chief interest. The application of Bayes’ theorem to the model in Section 4, yields:
@@
$(\hat{\bm{c}},\hat{\bm{\tau}},\hat{\theta})=\underset{\bm{c},\bm{\tau},\theta}{\arg\max}~{}p(\bm{c},\bm{\tau},\theta\,|\,\bm{x},\bm{y}).$ (5)
The point estimate, $(\hat{\bm{c}},\hat{\bm{\tau}},\hat{\theta})$, is of the mode of the parameters’ posterior density. Next, we discuss how to solve this problem using a stochastic gradient ascent method. Our procedure follows standard practices in deep learning [35], with some adaptations to our multi-unit soft sensor model.

### v3

# Multi-task and few-shot learning in virtual flow metering

@@ Abstract
Recent literature has explored various ways to improve soft sensors by utilizing learning algorithms with transferability. A performance gain is generally attained when knowledge is transferred among strongly related soft sensor learning tasks. One setting where it is reasonable to expect strongly related tasks, is when learning soft sensors for separate process units that are of the same type. Applying methods that exploit transferability in this setting leads to what we call multi-unit soft sensing.
This paper formulates a probabilistic, hierarchical model for multi-unit soft sensing. The model is implemented using a deep neural network. The proposed learning method is studied empirically on a large-scale industrial case by developing virtual flow meters (a type of soft sensor) for 80 petroleum wells. We investigate how the model generalizes with the number of wells/units. We demonstrate that multi-unit models learned from data from many wells permit few-shot learning of virtual flow meters for new wells. Surprisingly, regarding the difficulty of the tasks, few-shot learning on 1-3 data points often leads to high performance on new wells.
keywords:
soft sensor, hierarchical model, neural network, multi-task learning, few-shot learning
@@ 1.1 Soft sensing
In essence, a soft sensor is a mathematical model that operates on available process measurements ($x$) to infer a variable of interest ($y$). Arguably, the most common form of a soft sensor is $\mathrm{E}\mathopen{}\left[y\,|\,x,\theta\right]\mathclose{}=F(x;\theta)$, where the conditional expected value of $y$ is modeled by a function $F$ with arguments $x$ and parameters $\theta$. Two important assumptions in the development of a soft sensor are i) that the process measurements, $x$, are informative of $y$, and ii) that the process measurements arrive with a higher frequency than measurements of $y$. Under these assumptions, a soft sensor can be used to monitor $y$ indirectly at times when only $x$ is measured; when $x$ is available online, a soft sensor can provide online predictions of $y$. Furthermore, when $y$ is a key performance or quality indicator, the process can be optimized based on the inferred values.
A soft sensor can be classified as being either model-driven or data-driven, based on the specification of $F$ [2]. In a model-driven soft sensor, the model $F$ is derived from first-principles. For example, $F$ may be a mechanistic model whose parameters represent physical properties of the process being modeled. Note that while most sensors rely on some kind of model to map the variable of interest to a physical quantity which is practical to process and record (e.g. an electric voltage), the term “model" here refers models which are not integrated and shipped as part of the sensor.
In a data-driven soft sensor, $F$ is a generic function learned from data, e.g. a linear regression model or a neural network. Soft sensors that combine the model-driven and data-driven approaches are sometimes referred to as hybrid soft sensors. Regardless of the soft sensor type, the model should be calibrated to available measurement pairs $\mathcal{D}=\{(x,y)\}$ before being used for prediction. Commonly used methods include least-squares or maximum likelihood estimation (MLE). When a prior is available, maximum a posteriori (MAP) estimation or Bayesian inference can be used to calibrate the model [9].
@@
The disadvantages of model-driven soft sensors have motivated much of the development of data-driven soft sensors. The attraction to data-driven soft sensing comes from its promise to simplify and reduce the cost of modeling by learning the model from historical data using statistical inference or machine learning techniques. A challenge with data-driven soft sensing is to learn a model from information-poor data, where the paucity of information may be due to a low data volume, frequency, variety, or quality [2]. The problem of low data variety in particular provides a challenge for data-driven methods, since these can generally not be expected to extrapolate outside of their training distribution. These data challenges are closely tied to the motivation for implementing a soft sensor, which is to infer a key variable that is measured infrequently. Put differently, a soft sensor is most valuable in information-poor environments, where one would expect data-driven methods to perform poorly. To achieve a good predictive performance in such environments, it is imperative to be data efficient. Model-driven soft sensors achieve this by having few parameters and strong priors. On the other hand, data-driven soft sensors, which may have more model parameters than observations to learn from, often rely on generic regularization techniques and weak priors. In effect, this means that their model is learned from scratch for each application. Thus, we argue that data efficiency is a pressing issue in data-driven soft sensing.
@@ 1.2 Multi-unit soft sensing
One strategy to making data-driven soft sensors more data efficient is to exploit transferability [10]. A general description of transferability is the ability of a learning algorithm to transfer/share knowledge between different learning tasks to generalize better. In soft sensing, transferability entails learning across multiple soft sensor tasks. If the tasks are strongly related (i.e., share important characteristics), one may expect transferability to help soft sensors generalize. Various ways to employ transferability in soft sensing are covered in the following, where we elaborate on how different learning methods have been used to transfer knowledge between different soft sensor tasks.
@@
A particularly relevant case for transferability is when the same type of soft sensor is to be developed for several similar, but physically different processes. An example is when a product quality indicator (soft sensor) is to be developed for several, almost identical production lines. In this case, it is reasonable to expect the different soft sensor tasks to be strongly related. We call the application of transferability in this setting multi-unit soft sensing, where unit refers to a physical process being modeled by a soft sensor. [^4] In other works, a unit may be referred to as an entity, subject, object, or realization for which we have multiple observations. In statistical terms, it would be called a unit of analysis with repeated measurements. Also note that a unit may represent a unit operation, which is a term used in chemical engineering to refer to a basic step in a larger process. For example, units may refer to a set of pumps, valves, solar panels, distillation columns, geothermal wells, or robots. The setting of multi-unit soft sensing is illustrated in Figure 1.
@@
Data-driven multi-unit soft sensing will be framed in mathematical terms in our problem statement. Interestingly, the framing allows us to draw a parallel to model-driven soft sensing. Suppose that we have access to a large multi-unit dataset, in terms of the number of observations and units. One could then expect that a learning algorithm with transferability will converge to a prototypical representation of the data (which we can view as a strong prior model). When facing new units, this prototypical representation can be leveraged to quickly learn a soft sensor from a few data points. In machine learning terminology, this would be an example of few-shot learning, where a model performs well on a task after being calibrated to a few (often less than ten) data points [11, 12]. The parallel to model-driven soft sensing is that a first-principles model and a learned prototypical model take the same role of representing the general process.
@@
Figure 1: Illustration of multi-unit data generation. Here, each unit $F_{i}$ corresponds to a physical unit performing the same basic operation. The $M$ units are similar to each other, and can be thought of as different realizations of some “prototype" unit, denoted here by $F$. Still, each unit also has its own unique physical characteristics, resulting in datasets $\mathcal{D}_{i}$ with different distributions.
@@ 1.3 Virtual flow metering
Data-driven VFMs attempt to address the cost issue with first principles VFMs by lowering the model setup and maintenance costs. In addition to reducing manual labor related to physical modeling, data-driven VFMs typically also reduce the need for manual parameter tuning by allowing for automatic model tuning in the presence of new data [19, 20].
If the flow conditions are sufficiently predictable, data-driven models based on simple first principles equations (which need to be simple enough to allow for automatic estimation of all model parameters from data) may yield satisfactory prediction accuracy [21]. If prior knowledge about flow conditions is lacking, it may be possible to compensate by learning from more data; in such settings, steady-state VFMs based on deep learning have shown promising results [22, 23, 24, 13]. In the presence of data with high time resolution, methods based on deep neural networks have also been shown capable to model dynamical flow phenomena like slugging and transient behavior following step changes [25].
It is commonly believed that the successful use of data-driven VFMs based on deep neural networks requires large amounts of data [13]. Such a requirement would limit their applicability, since the wells with the worst instrumentation (and hence, the least data) are typically also the wells were VFMs are needed the most. Thus, the application of data-driven VFM to wells with few observations is a problem of high interest.
@@ 1.5 Paper outline
The paper proceeds by discussing related works , before giving a mathematical problem formulation. The multi-unit soft sensor model is then described, followed by the derivation of the learning method. An empirical study based on a large-scale industrial dataset is presented, before concluding remarks are given. Some implementation details and additional results are gathered in an appendix.
@@ 1.6 Notation
The normal distribution is denoted by $\mathcal{N}(\mu,\Sigma)$, where $\mu$ is a mean vector and $\Sigma$ is a covariance matrix. If $X$ is a normally distributed random variable, we denote a sample by $x\sim\mathcal{N}(x\,|\,\mu,\Sigma)$. Generally, we use the letter $p$ to denote a probability distribution, $p(x)$, or a conditional probability distribution, p(y|x)p(y\,|\,x). We sometimes use the letter $q$ for probability distributions that are approximative.
@@ 2 Related works
Finally, knowledge can be transferred between soft sensors for similar, but physically different processes. This setting is referred to as cross-entity transfer learning in [10]. The idea behind cross-entity transfer learning, is that, if two entities/processes/units are physically similar, data from one unit will be informative to the operation of the other unit. Thus, with data from multiple units, soft sensor performance may be improved by pooling the data and learning across units. This learning paradigm has high potential since many soft sensors are developed for processes that are commonly found in the industry; consider a soft sensor for a specific pump unit, which may be of a model produced in large numbers and installed in many process systems. Despite the high potential, few examples of cross-entity transfer learning can be found in the literature (one reason may be that multi-unit datasets are rarely shared). One example is given in [23], where a data-driven virtual flow meter is learned from data from many oil wells using multi-task learning.
2.1 Nomenclature
In our work, we prefer to use the term multi-unit soft sensing since it reflects that learning happens across units without specifying by which means (e.g. transfer learning or multi-task learning). In settings where the means of transferability is relevant, we will however state which paradigm best describes the learning problem in question. Briefly summarized, pretraining will be done through the use of multi-task learning while few-shot learning will be done through a finetuning/calibration strategy which constitutes a form of transfer learning. We will use the term few-shot transfer learning in settings where we believe this deserves emphasis. In the experimental case study we will apply cross-entity transfer learning; the terms unit and task will be used interchangeably in this particular context.
@@ 3 Problem statement
We consider the underlying process $p(x,y)$ that generates data for any unit in a population. Let $c\in\mathbb{R}^{K}$ be a latent context variable with distribution $p(c)$. The context identifies a unit in the population supported by $p(c)$. We can then think of the process as the marginalization over all units, i.e., $p(x,y)=\int p(x,y\,|\,c)p(c)dc$. With this model, $p(x,y\,|\,c_{i})=p_{i}(x,y)$, for a given realization $c=c_{i}$. This implies that $p(y\,|\,x,c_{i})=p_{i}(y\,|\,x)$, showing that we can model the behavior of all units by a single conditional distribution.
We now state that the goal of multi-unit soft sensing is to learn from $\mathcal{D}$ a distribution q(y|x,c)q(y\,|\,x,c) that approximates p(y|x,c)p(y\,|\,x,c). We contrast this to the single-unit soft sensing, where $p_{i}(y\,|\,x)$ is approximated using $\mathcal{D}_{i}$, for $i=1,\ldots,M$; that is, where $M$ soft sensors are learned separately.
Our work aims to study empirically and theoretically the predictive performance of the approximation q(y|x,c)q(y\,|\,x,c). First, we want to investigate how performance scales with the number of units $M$. Second, we want to examine the few-shot learning performance of q(y|x,c)q(y\,|\,x,c) as the context $c$ is calibrated to few observations of a new unit (whose data was not used to learn $q$).
@@ 4 Model description
We consider a hierarchical model of the observed target variables for a set of units. The model has three levels, as illustrated graphically in Figure 2. The top-level has universal parameters ($\theta$) that are shared by all units and observations. The middle unit layer has parameters ($c$ and $\tau$) that capture the specificity of units. At the inner observation level, the target variable ($y$) is explained by the explanatory variables ($x$) and the context ($c$) of the observed unit. The target observation is subject to a measurement error ($\varepsilon$) with a unit-specific variance ($1/\tau$). The relationship between these variables is modeled using the universal parameters ($\theta$). The hierarchical structure allows the model to capture variations at the level of units and observations. A mathematical description of the model is given next.
@@
where $\varepsilon_{ij}$ is the error term. The error, here interpreted as measurement noise, is modeled as $\varepsilon_{ij}\sim\mathcal{N}(0,1/\tau_{i})$, where the precision (or variance $\sigma_{i}^{2}=1/\tau_{i}$) is allowed to differ between units. The function $f$ is modeled by a neural network with parameters $(c,\theta)$ and arguments $x$. The specific architecture of $f$ is described in the appendix.
@@
The model in (1) results from several simplifying assumptions: 1) the dimension of the context variables, $K$, is fixed and considered a design choice; 2) the model is conditional on $x$, and thus not a generative model of the observed data; 3) the noise is Gaussian; 4) the model is homoscedastic since the measurement noise variance, $\sigma_{i}^{2}$, is fixed (per unit). Assumption 2 leads to a discriminative model which is suitable when the soft sensor will be used for on-line prediction. Assumption 3 could be relaxed (e.g. by generalization to exponential family) without much technical complication, but is kept for reasons of simplicity. Relaxing assumption 4, for example by allowing $\sigma_{i}$ to vary with $x$, would give a more complex heteroscedastic model, which may be harder to learn in a data-scarce setting.
@@ 4.2 Likelihood
$\ell(\bm{c},\bm{\tau},\theta\,|\,\mathcal{D})$= $\log p(\bm{y}\,|\,\bm{x},\bm{c},\bm{\tau},\theta)$ (3)
@@
We use the short-form notation $\ell(\bm{c},\bm{\tau},\theta\,|\,\mathcal{D})$ for the log-likelihood when it is considered as a function of the parameters ($\bm{c}$, $\bm{\tau}$, $\theta$), given data $\mathcal{D}=(\bm{x},\bm{y})$. We will show shortly how to perform MAP estimation of the parameters, given the log-likelihood in (3) and log-prior in (2).
@@ 5 Learning method
In model learning, the posterior density of the parameters $(\bm{c},\bm{\tau},\theta)$ is of chief interest. The application of Bayes’ theorem yields:
@@
$(\hat{\bm{c}},\hat{\bm{\tau}},\hat{\theta})=\underset{\bm{c},\bm{\tau},\theta}{\arg\max}\leavevmode\nobreak\ p(\bm{c},\bm{\tau},\theta\,|\,\bm{x},\bm{y}).$ (5)
[…]